Deduce the dimension of planck’s constant . Which other quantity has same dimension?

Grade:11

1 Answers

Akarsh Tripathi
6 years ago
The dimension of Plank`s contant can be obtained from any of the equations in which it appears, e.g. h = E/ν, λ = h/p, etc.The Planck`s constant and angular momentum have the same dimensions.i.e. [h] = [M L^2 T −1] and [l] = [r] [p] = [L] [M L T −1] = [M L^2 T −1]